The demand function for Product X is given by: Qdx = 10 + 0.06I - 2Px - 0.5Py + 0.7Pz  where Px  Price of good X $9.00 

Py Price of related good Y  $4.00

Pz Price of related good Z $10.00 

I Income  $250.00 

a. (i) Calculate the own Price elasticity of demand (PED) for Good X.                           

                           


1
Expert's answer
2021-10-24T18:09:58-0400

Elasticity of demand depicts how much consumer responds with the change in the price level. 


own ped , "Ed_x=\frac{Qd_x}{dP_x}\times\frac{P_x}{Qd_x}"


"=-2\times \frac{9}{10+0.06(250)-2(9)-0.5(4)+0.7(10)}\\=-1.5"


Therefore, |Edx1.5>1, elastic